11. Following are three sample equations. Plot them on a graph in which Q is on the vertical




axis and P is on the horizontal axis. Then transform these equations so P is expressed in




terms of Q and plot these transformed equations on a graph in which P is on the vertical




axis and Q is on the horizontal axis.




a. Q = 250 - 10P




b. Q = 1,300 - 140P




c. Q = 45 - 0.5P

1
Expert's answer
2022-03-08T12:18:09-0500

Plot them on a graph in which Q is on the vertical axis and P is on the horizontal axis.


When P =1, 2, 3, 4, and 5; Q is:




The graph is:



Then transform these equations so P is expressed in terms of Q and plot these transformed equations on a graph in which P is on the vertical axis and Q is on the horizontal axis.


a)

Q = 250 - 10P


P = "\frac{250 - Q}{10}"


"P = 25 - 0.1Q"


b)

Q = 1,300 - 140P


P  = "\frac{1300 - Q}{140}"


"P = 9.2857 - 0.0071Q"

      

c)

Q = 45 - 0.5P


P = "\frac{45 - Q}{0.5}"


"P = 90 - 2Q"


Plot graph

When Q =1, 2, 3, 4, and 5, P is:


The graph is: